OpenStudy (anonymous):

how can i find the inverse of y=e^(2x+1)

OpenStudy (kainui):

Swap the y and x values and solve for y.

OpenStudy (anonymous):

i understand that part but how yould i solve? would i take ln of e?

OpenStudy (anonymous):

Here you go.

OpenStudy (anonymous):

thank you. did you take ln of both sides to cancel out the e?

OpenStudy (aum):

Yes. \(\large \ln(x) = \ln(e^{2y+1}) = 2y+1\)

OpenStudy (anonymous):

thank you so very much!

OpenStudy (aum):

yw.